OpenStudy (anonymous):

find the solution of 5c - c^2=0 a) 0 & 5 b) 0 & -5 c) 0 , 2 , & 5 please help !

OpenStudy (jack1):

5c - c^2=0 so factorise 5c - c^2=0 c(5-c) = 0 therefore either c = 0 OR 5-c = 0 5 = c so 0 is a solution, and 5 is a solution

OpenStudy (anonymous):

thank you so much ! :)

OpenStudy (jack1):

sall good, make sense?

OpenStudy (anonymous):

yeah , it's pretty easy